Child pages
  • 201702 - Public report

Versions Compared

Key

  • This line was added.
  • This line was removed.
  • Formatting was changed.

Excerpt
div

Unscheduled:

  • February 22nd: The link between one of the routers in MIUN and SUNET flapped due to a misconfiguration during a troubleshooting session.
  • February 18th: Lost contact with the following hosts located on the IaaS:

    ds-legacy.sunet.se
    ds-test.swamid.se
    registry-test.swamid.se
    satosa-proxy-2.swamid.se
  • February 17th: One of the links towards SMHI was down. It was due to a fiber problem at the provider's network. The problem is now solved and the connection is up again.
  • February 16th: We experienced a problem with the DNS for SUNET. This was caused by reconfiguration of dns zone sunet.se losing entries and the long restoration period (25min after detection) was due to the hard handled CVS revisioning system.
  • February 15th: Ther was a routing issue at SUNET provoked by a hardware failure on one uplink to NORDUnet. The detailed RFO can be followed up on NORDUNETTICKET-4264.
  • February 12th: We experienced sporadic packet losses toward multiple hosts in IAAS. Provider reconfigured performance parameters and solved the performance issues.
  • February 7th: Host v-mailfilter4.sunet.se is down. Provider has restarted the server on their end and freed some space. The problem was due to a known issue that should be solved at the next rollout.
  • February 7th: We lost power redundancy on both juniper routers located at HKR: hkr-r1.sunet.se and hkr-r2.sunet.se. This was due to a power outage on site. The power is restored and all connections are up and running.
  • February 4th: After a planned power maintenance at Södertörn Högskola they experienced problems with their core routers. Customer fixed it by installing new component.
  • February 1st: An emergency routing switchover was performed on router gbg7-r1.
  • January 29th: A FPC card in ltu-r1 was faulty and was replaced.

Scheduled:

  • February 28th: The router in Filipstad was upgraded to a new software release.
  • February 28th: We replaced our RAMAN in Junosuando (KIR4) to fix a hardware issue.
  • February 27th: We restarted FPC card in one of our routers in Stockholm.
  • February 26th: Supplier performed a planned maintenance in their network on the connection Kiruna - Pajala.
  • February 21st: 8 PSM card is installed in HIG2.
  • February 20th: We had to fix a DCP error on cth-r1 by reseating FPC0
  • February 16th: The hostname and loopback IP on m1fls have been successfully changed.
  • February 7th: The server e-mailfilter03 is upgraded with new hardware.
  • February 5th: Fiber provider performed a maintenance in the network between Åkroken and Älvsbyn.
  • February 2nd: Fiber provider performed a planned maintenance between Hedenlunda and Norrköping.
  • February 1st: NPE performed longterm OTDR measurement between KOB1 and HDA. Found out that the ODF in HDA was bad. T2 and NPE swapped to a new pair.

Scope

Unscheduled

Scheduled

Total

Hardware

1 4 pcs 1d 1114:0131

2 pcs 00:2127

3 pcs 1d 1115:2208

Link

1 2 pcs 0020:079

3 4 pcs 091d 08:4126

4 6 pcs 091d 28:4835

None

3 0 pcs 0100:00

5 1 pcs 2300:2510

8 1 pcs 1d 00:2510

Routing

0 2 pcs 00:30

0 1 pcs 00:21

0 2 pcs 00:51

Software

1 pcs 00:2430

0 1 pcs 11:7

1 2 pcs 0011:2437